Observing Run Synopses


Commissioning run: 16-19 July 2001 (2+2)

This run was postponed from 5-6,16-18 April 2001 because of delays in instrument delivery.

Although there were some minor equipment problems, overall the PN.S turned out to run quite smoothly and according to expectations. The measured telescope-to-detector efficiency of 33.3%±1.0% was in accord with the expected 35.0% (see here for the derivation).

The weather was less cooperative: dust, clouds, poor seeing. We ended up with relatively little high-quality data. An overview of the targets observed and data quality can be found here.

Photos from the run can be found here and science images here.


Run #2: 13-18 September 2001 (3+3)

The observing conditions were mostly good, and the instrument performed well. We observed several galaxies, getting especially good data on NGC 821, NGC 7332, and NGC 7457.

Photos from the run can be found here and science images here.


Run #3: 7-14 March 2002 (4+4)

Poor observing conditions (poor seeing, cloud cover, ice storm, etc.) meant that we got effectively one night of good data in eight nights' observing. Much better than nothing! We got about a half night of data on each of NGC 3379 and NGC 4494.

Summary:
NightHours on Primary TargetsSeeing1" HoursOfficial time lostDark timeGalaxies
1 (7 Mar)6.31.9"1.70.09.6N3379, N4494, M51
2 (8 Mar)2.82.0"0.70.09.6N3379, N4494, M94, N2683
3 (9 Mar)6.81.5"3.20.09.6N3379, N4494, N2549
4 (10 Mar)0.02.4"0.05.89.6M94
5 (11 Mar)0.02.6"0.05.89.6M94
6 (12 Mar)0.00.09.59.5
7 (13 Mar)1.42.3"0.35.09.5N3379, N4494
8 (14 Mar)6.81.3"3.70.09.4N3379, N4494, N2683
Total24.11.6"9.627.576.3

First 4 nights were UK; second 4 were NL.
Official dark time lost: 27.5h out of 76.3h (5.8 UK, 21.7 NL).
Estimated 1" time acquired: 9.6h (5.6h UK, 4.0 NL)

Photos from the run can be found here and science images here.


Run #4: 8-13 October 2002 (3+3)

H3 M31 run, Dave Carter, Helen Merrett, AJR, NGD and (first three nights) KK. HM/NGD/AJR had an overnight stop in Madrid, KK a day later.
We convinced the observatory to allow us to try out the MARCONI2 detector on the right arm. This did not improve the high read-noise and cross-talk, which fact led to a thorugh investigation of the problem (which was also present when the detector was replaced by the EEV13). The technical team managed to get acceptable read-noise, but on the third day there was a breakthrough when the culprit was identified (controller power supply). We subsequently marked the preferred power supplies for future use.
Mixed weather for 8/10 and 9/10 with the dome at times closed due to rampant humidity - the seeing was also quite poor. During 10-13/10 the weather was good with low and stable cloud patterns and the seeing steadily improved. However on the last three days approximately one hour per night was spent on a GRB override.
Some 130 (?) fields were observed with a standard exposure of 900s, the results are promising.
Photos from the run can be found here.


Run #5: 28 February - 7 March 2003 (4+4)

Primary programme. EEV12+Marconi2 requested. Marconi2 is cosmetically better and has slightly higher QE. Observers AJR, HM, NGD and NN.

AJR, HM and NGD arrived early in order to investigate "clunk", suspicion is that the grating unit might be subject to occasional bimodal disalignment. On Feb 27 we dismantled it and convined ourselves that the pushers which press the grating against the mid-line might be too weak, allowing lateral movement and maybe rotation within the available space (some mm) - see, however, John Hart's email of 1/3/03. Trial: K. Dee added a second spring, and we inserted a sawn-off M4 nylon screw head into the hole, further increasing the pressure. Blew dust off. Possible resulting changes in grating alignment indicated by PNSGANG=26.6-2 in headers.

We experienced great difficulty in getting Marconi2 focussed (right arm) and optimum result was 3.0 (c.f. 2.5 for EEVs). Currently not understood. Gives rise to different optimum telescope focus on L/R arms. During first night the network went down, causing some difficulties with target optimalisation. However no technical problems otherwise and good data obtained. Confirmed that Marconi2 is some 10% higher in QE, though thicker and thus more prone to cosmic ray events.

On days 2 through 6 we likewise stuggled with focus, investigating possible culprits (tilted filter, the modifications to the filter unit). Cameras not parfocal, as indicated by telescope focus runs. TCS failed at the very end of night 2, sky flats were made with WHT at zenith. Night 3 had seeing of ~3".

Night 4, reverted to EEV13 on right arm. Seeing was again ~2-3".

Night 5 thru 8 had clouds off and on but reasonable seeing.

Summary:
NightHours on Primary TargetsSeeing1" HoursOfficial time lostDark timeGalaxies
1 (28 Feb)6.51.5"2.80.09.8N3379, N4339, N4494
2 (1 Mar)6.51.5"3.00.09.8N2549, N3379, N4339, N4494
3 (2 Mar)0.51.7"0.20.09.8N3384, N4494
4 (3 Mar)1.22.1"0.30.09.8N3384, N3377, N4494
5 (4 Mar)6.61.3"4.10.09.7N2549, N3379, N4339, N4494
6 (5 Mar)0.01.2"0.03.09.7N2549, N3379, N4374
7 (6 Mar)0.01.2"0.04.59.7(N2549), (N3115), N4374
8 (7 Mar)0.00.09.09.7(N3379), (N2549), (N4374), (N3384)
All21.31.4"10.416.578.0

First 4 nights were NL; second 4 were UK.
Official dark time lost: 16.5h out of 78.0h (all from UK).
Estimated 1" time acquired: 10.4h (6.3h NL, 4.1h UK)

Photos from the run can be found here.


H3 JKT-imaging-Run #1: 1-5 June 2003 (5NL)

Mirela Obric, PhD student at Kapteyn, obtained photometric-quality images of NGC7013 and NGC5866 with the Harris-R filter (6373/1491) for use in our modelling of these S0 galaxies.
5 good nights, according to logs.


H3 INT-wfc-Run #1: 16-21 August 2003 (6NL)

M31 survey run using the wide-field-camera at INT. Obsevers Dave Carter and HM.

Good seeing on 4 nights (av 1.2"), cloud cover for 1 night. 23 fields imaged in M31's halo.


Run #6: 29 Sept - 5 Oct 2003 (4UK+3NL)

M31 survey extension - major axis, stream and northern spur. Observers HM, AR, NN, DC.

3.5 clear nights. Varied seeing (1.0-2.0"). 88 fields observed.


Run #7: 12- 21 april 2004 (6UK+4NL) (P39 Mer + N12 Dou)
(12-15 Apr NL time; 16-21 Apr UK time)
		arrives		leaves
Koen		11/4		17/4
Aaron		12/4		21/4    (arrives La Palma 12/4 10:15)
Nigel           15/4            18/4    (arrives La Palma 15/4 15:30)
Magda		17/4		23/4	(arrives La Palma 16/4 20:45)
Ortwin		17/4            23/4    (arrives La Palma 16/4 20:45)

Ice, snow, clouds for 5.5 out of first 6 nights. Last 4 nights clear with reasonably good seeing.
Extensive practise with swapping filters while observing.
Galaxies NGCs 4564, 4552, 5846 observed.

Summary:
NightHours on Primary TargetsSeeing1" HoursOfficial time lostDark timeGalaxies
1 (12 Apr)0.00.08.58.5
2 (13 Apr)0.00.08.68.6
3 (14 Apr)0.00.08.58.5
4 (15 Apr)3.11.5"1.43.88.4N4564, N5846
5 (16 Apr)0.00.08.48.4
6 (17 Apr)0.12.3"0.07.68.3N4564
7 (18 Apr)5.01.3"2.80.08.3N4564, N5846
8 (19 Apr)5.51.2"3.70.08.3N4564, N5846
9 (20 Apr)5.01.3"3.10.08.2N4552, N5846
10 (21 Apr)3.01.0"2.90.08.2N4552, N4697, N5846
Total21.71.25"13.945.483.7



Run #8: 1- 10 april 2005 (5UK+5NL) (P37 Mer + N14 Dou)
(1-5 Apr NL time; 6-10 Apr UK time)
First 4 nights with great seeing. Then a couple of nights with poor seeing and some cloud interference. Then a couple of decent nights. Then clouds moved in again.

Good seeing on DIMM often corresponded to awful seeing at telescope. More detailed information on this matter is found with email to Neil O'Mahoney, data table (original seeing table) and fig 1, fig 2 and fig 3. Here is Chris Benn's reply. The problem seems to have largely been a high mirror-dome temperature differential, but this wouldn't explain good seeing on the guider at times during poor PN.S seeing. Let's keep an eye out for this phenomenon again and document it. Here is Chris' further investigation into the seeing issue.

Galaxies NGCs 3377, 3608, 4374, 5846 observed.

Summary:
NightHours on Primary TargetsSeeing1" Hours (prim targ)Official time lostDark timeGalaxies
1 (1 Apr)6.01.1"4.10.08.9N3377, N4374, N5907
2 (2 Apr)8.01.1"6.30.08.8N3377, N4374, N5846, pimdouglas
3 (3 Apr)8.51.1"7.20.08.8N3377, N4374, N5846
4 (4 Apr)8.00.9"9.00.38.8N3377, N4374, N5846
5 (5 Apr)6.52.2"1.50.08.7N3377, N4278, N5866
6 (6 Apr)2.02.0"0.40.08.7N3377, N3384, N5363, N5866
7 (7 Apr)5.01.5"2.22.78.7N3608, N4473, N5846, pimdouglas
8 (8 Apr)5.71.7"2.00.08.6N3608, N4473, N5363
9 (9 Apr)1.01.8"0.30.08.6N2683, N3377, N3384, N5866
10 (10 Apr)3.51.9"2.63.08.6N3608, N4564, N5866
Total54.21.6"35.66.087.2



Run #9: 27- 30 march 2006 (4UK only) (05A P37 Mer)

Night 1

Edo Noordermeer, Nigel Douglas, Lodovico Coccato present at start. T.O. Juan Carlos.
PNS cupboard in good shape, key located in Ops Key cupboard. Confirmed filters A, AB, C in box 3, each in wooden box; filter B (mounted in PNS at start of the run) needs wooden box. Confirmed multiple "special tools" for exchanging filters, though only one of the three is magnetic. Set up PC: special cable now appears to be permanently connected underneath floor to break-out boxes. Needed to add doskey.com to the PC setup. On arrival, lots of cirrus and thicker clouds with breaks. Plan to start with filter B targets throughout night1. During installation the PNS cables A and B had been connected to the wrong connectors at the connectors at Cass. Should be 4/1 respectively.
Also, the filter was left at B6 on the previous run, not B0 as indicated on the note attached to the PNS. We used the turning tool to confirm it to be B0 and this agreed with our spectra. We forgot to update the GUI (still at B6), but fixed the header manually the next night. Got underway with no major problems, seeing 1.3/1.1 (L/R).
DIMM seeing is 0.8, we measure 1.2 on 1800s images


Night 2

Good news in the afternoon, read noise on EEV? reduced. Technical problems at start of observing. DAS showed wrong directories, which was confused by occasional absence of image. Other images present but no data or constant value (eg 512). Problem traced by Don Carlo via phone, is miscommunication between CCD-controller and DAS. Not reproducable and thus not fixable. Problem did not re-occur, so no follow-up action undertaken.
Mediocre seeing (1.5 - 1.7") throughout the night. Note that DIMM seeing is systematically lower again (~1.1").
ND leaves La Palma at end of the night.



Night 3

Switched from filter B to AB in the afternoon. Focussing of spectrograph turned out very difficult. Especially on left arm, no perfect focus position could be found. Settled at optimal value (~3.2 pixels FWHM). Given the poor seeing in the previous nights, slightly non-optimal focus is not expected to influence science results.
Managed to obtain good data throughout the night. The seeing gradually improved from about 1.6" at the beginning of the night to about 1.1" at the end. This was, however, accompanied by the approach of a large front of clouds, visible on the weather satellite images. Poor conditions are expected for the next night.



Night 4

Evening started with lots of clouds overhead, but weather cleared up in the first few hours. Cirrus present throughout the entire night though, so not photometric. Seeing similar as previous nights: 1.5" - 2.0".
At 5.15h, Taurus (controlling the DAS) crashed and needed to be rebooted. Lost about 45 minutes as a result. At 6.00h, system up and running again, so could still take 2 exposures on N 5866.


End of run

PN.S dismounted, equipment stored in PN.S cupboard. Instrument contains filter AB, filter B stored in box 4 (round plastic container).


Summary:
Night Hours on Primary TargetsSeeing1" Hours (prim targ)Official time lostDark timeGalaxies
1 (27 Mar) 6.0 1.3" 3.6 0:00h 9:00h N3608, N4473, N4742, N5846
1 (28 Mar) 8.0 1.7" 2.8 0:00h 9:00h N3608, N4473, N5846
1 (29 Mar) 8.0 1.3" 4.7 0:15h 9:00h N3115, N4278, N4374, N5866
1 (30 Mar) 7.75 1.7" 2.7 0:45h 9:00h N3115, N4278, N4374, N5866
Total 29.75 1.5" 13.2 1:00h 36:00h



Score to date:

UK hrsNL hrsTotal hrs
Dark hours acquired:170.1155.1325.2
Official hours lost44.051.495.4
Hours on target:58.363.0121.3
1" equivalent hours:29.739.869.5





Run #10: october 06

The "advance party" arrived to set up the H-Alpha camera (HaC). Mike and Nigel were on mountain on October 18, Lodo on 20th. In addition we had asked a software specialist, Erik de Wit to join us to fine-tune the new control system and interface for the PN.S. This was needed because of the extra requirements of the PN.S + HaC. The controller used since 2001 can operate two motors (shutter and mask) and can be addressed via a parallel port. Controller II will be required to operate four motors since the HaC has a wide and a narrow-band filter which can be exchanged by motor drive, making four in total. In addition Controller II is designed to be more flexible, with the possibility to send commands to other devices via its own serial port. This is needed for the HaC iris shutter which has a driver which responds to serial commands, and Controller II is capable of instructing four such devices. MM and ND unpacked the boxes from the RSAA and were pleasantly surprised to find the HaC to be pre-assembled (rather than in parts as expected). The consignment also contained Controller II which had been tested by ND in Australia and left there to be freighted along with the other hardware. The wide-band and narrow-band filters (the latter rejected) were also present. The replacement narrow-band filter had arrived independently and was stored in the PN.S cupboard. The filters have not been inspected or tested by us. We encountered the following complications and/or setbacks.
  • The HaC was delivered without a flange on which to mount a CCD cryostat, and without documentation on how to build one (i.e. the position of the focal plane was not marked and nor could it be easily derived from the documentation provided).
  • The Controller had suffered minor damage in transit; screws were loose, a motor card was broken, etc.


Information was sought and quickly received from RSAA, allowing the ING to start building a cryostat flange. The obvious problems with the Controller were repaired but there were still malfunctions: To the observers on this run we request that you look for evidence, in the OIII observations, of any extra flexure. We also request some careful standard star observations under photometric conditions in order to check the losses introduced by the dichroic.
Important note on OIII filter changes:
What used to be the square access flange is now one side of the dichroic holder. Remove the OUTER 8 screws and using the handles to pull the unit out as well as providing support against it falling, remove the dichroic holder. Have a safe place ready to place the dichroic. Exchange or rotate the filter as required then re-insert the dichroic and the 8 screws. We have a new piece of paper for the filter data to be filled in (should also be in the GUI of course) current filter AB-0o




Sunday, the day before the first night; focusing the spectrograph.

We encoutered a software problem during the PNS focusing procedure: the new upgrade of the CCD controller software looks for the shutters on the detectors. Since we have our own shutter and no shutters on the two PNS detectors, the software gets crazy. Result: it never stop the integration. We temporanely solved the problem (suggestion of the tecnicians) by taking dark exposures with our shutter open. In "dark-mode" the new software is not looking for the shutter and it will not complain. Our shutter is open and we get light anyway. We were therefore able to focus the instrument (right focus better than the right one).
A new logsheet is available now at the manual webpage




Night 1, Monday 23 Oct, 2006.


Halpha filter mounted: Halpha not available
Sky conditions:nearly photometric
PNS filter mounted: AB-0
Observers: LC, MM, EN


The Halpha CCD generates a lot of interferences with the two PND detectors. Moreover, it has a extremly bad read out noise. For this night the Halpha camera is therefore not available. Please have a look to the report at www.astro.rug.nl/~coccato/Halpha.html


Night 2, Tuesday 24 Oct, 2006.


Halpha filter mounted: Halpha not available
Sky contitions: rain first part, high humidity second part
PNS filter mounted: AB-0
Observers: LC, EN


There are still problems with the Halpha CCD R.O.N..

Dome closed for rain and humidity

Night 3, Wednesday 25 Oct, 2006.


Halpha filter mounted: Halpha not available
Sky contitions: clouds and hig humidity all night
PNS filter mounted: AB-0
Observers: LC, EN


There are still problems with the Halpha CCD R.O.N..

Night 4, Thursday 26 Oct, 2006.


Halpha filter mounted: Halpha not available
Sky contitions: high humidity first part,bad seeing second part
PNS filter mounted: AB-0
Observers: LC, EN

Night 5, Thursday 27 Oct, 2006.


Halpha filter mounted: Halpha not available
Sky contitions: high humidity first part,~1"5 seeing second part
PNS filter mounted: AB-0
Observers: LC, EN

Night 6, Thursday 28 Oct, 2006.


Halpha filter mounted: Halpha not available
Sky contitions: high humidity
PNS filter mounted: AB-0
Observers: LC, EN


RUN 3 nights NL, 3 nights NL
Observers MM, LC, KK, ND, NN Doing fine here. We're observing, there is some cloud from the tail end of a massive cloud system over Spain. Let's hope we escape the worst. Getting the instrument set up was simple - focusing took a couple of tweaks only this time. H alpha is not up, but we are working with the two OIII arms just fine. There was an interesting half hour chasing all cable connections until we discovered a second (now thoroughly taped over!!) parallel port on the PC. We're 3 (Mike is here too, very offended nobody is wishing him luck!), so that is OK for the car. Mike and Lodo are going down on the 17th (Sat), so the only crunch might be the 16th in the evening. No need to rent a car for that I'd say. With Nicola that will make 3 of us again. Weather bad at times, dome closed for humidity and/or wind on occasions. Seeing also variable, 1.2" at start, gradually increasing, up to >2". Briefly <1" on Feb 21. Other technical issues:
  • at times the PNS-console reports error messages which do not seem to have happened.
  • at times the DAS had to be reset
  • we noticed that sometimes the L images appeared to be tilted with respect to the focal plane - there was a gradient in the apparent seeing. But no associated gradient in the mask images.
  • WARNING: a reare but disastrous error
    ING afternoon-team changed the configuration file to ISIS, giving us wrong PA and other parameters. Effect was one good-seeing image that can be scrapped. So: always take a test image.
  • Preliminary throughput meas made of dichroic (see NN).
Other issues:
  • Paper I submitted (18/02, 2:40 UT) and accepted (21/02, 2:54 UT) during this run!
  • Wooden box reqd for B filter, this will remove the remaining ambiguity of where filters live.
  • WARNING: a reare but disastrous error
    ING afternoon-team changed the configuration file to ISIS, giving us wrong PA and other parameters. Effect was one good-seeing image that can be scrapped. So: always take a test image.

Summary of the run: ??
Night Hours on Primary TargetsSeeing1" Hours (prim targ)Official time lostDark timeGalaxies
1 (23 Oct) 8:00 1"0 8.0 2:00 10h N7013, N1023(E+W), N2681
2 (24 Oct) 0:00 na 0.0 10:00 10h
3 (25 Oct) 0:00 na 0.0 10:00 10h
4 (26 Oct) 5:00 2"5 0.8 5:00 10h NGC 1023 E PA90; N2681
5 (27 Oct) 5:30 1"4 2.8 4:30 10h N1023(E+W), N2681
6 (28 Oct) 0:00 .. .. 10.00 10h
Total 18:30 11.6 41:30 60h



Run #11: 13-21 February 2007

First night Feb 13): Observers: Lodovico Coccato, Mike Merrifield, Konrad Kuijken
Clouds and bad (~1.5 - 2.0) seeing.
Second night (feb 14): Observers: Lodovico Coccato, Mike Merrifield, Konrad Kuijken
Veils and bad (~1.5 - 2.0) seeing.
Third night: (feb 15) Observers: Lodovico Coccato, Mike Merrifield, Konrad Kuijken
Clear sky and very bad (>2.0) seeing.
(note: plane hi-jacked and GC temporarily closed)
Fourth night: (16) Observers: Lodovico Coccato, Mike Merrifield, Konrad Kuijken, Nigel Douglas
Autoguider erratic. (Seeing 1.4-1.6) Clear sky and very bad (>2.0) seeing.
Fifth night: (feb 17) Observers: Nicola Napolitano, Konrad Kuijken, Nigel Douglas
Cirrus (seeing (1.8-2.2, 3 at end!) seeing.
Sixth night: (feb 18) Observers: Nigel Douglas, Konrad Kuijken, Nicola Napolitano
Good (1-1.3) seeing.
Worked out offset for C-0 filter (B-> C offset +0.5) Seventh night: (feb 19) Observers: Nigel Douglas, Konrad Kuijken, Nicola Napolitano
Dicroic throughput tests with the internal tungsten lamp. Prelim analysis shows throughput is 95%-97%.
During a period of humidity we took a set of dome flats at PA 0, 90, 180, 270. Eigth night: (feb 20) Observers: Konrad Kuijken, Nicola Napolitano, Nigel Douglas
Bad (2-3) seeing. Switched to a local group backup program. Search for PNe in the newly discovered SDSS dwarf galaxies: got one?
Ninth night: (feb 21) Observers: Konrad Kuijken, Nigel Douglas, Nicola Napolitano Good (subsecond-1.2) seeing. Discussion about the emission line in the dwarf being a PN. Prior and posterior probabilities abound. Summary of the run:
Galaxy nameFilterSeeing1" Hours
NGC 1161 B-0 xxx0.6
NGC 3608 B-0 xxx0.4
NGC 4473 B-0 xxx2.5
NGC 4473 C-0 xxx5.6
NGC 2768 B-0 xxx13.1
NGC 4564 B-6 xxx4.5
NGC 4742 B-0 xxx3.5
A few Lessons learnt "
  • Guiding: make sure to check the Autoguider focus (should be around 2000) and filter (AG-V).
  • At beginning of night ask T.O. to verify instrument configuration of CASS station. If it is set for another instrument the meaning of PA zero will change!
  • We had several instances when one of the CCDs returned completely blank images. This requires a reset of its DAS computer: at SYS prompt type dasreset pnsl for example.
  • Filter focus offsets:we determined that when changing from the C filter to B you need to apply a -0.5 turn of the capstans to refocus. Filter AB has the same offset from B, as does filter A. So it appears that A, AB and C are parfocal, with B half a turn offset, in the sense ANY to B: -0.5 turns and B to any +0.5 turns. Tilting filter B from 0 to 6 degrees does not alter the focus.
Software updates: The arcs and flats scripts were changed: they now switch on the comparison lamps automatically, and take care of the proper amount of neutral density.
If you start up IRAF in the pns subdirectory of /home/whtobs, you get two IRAF tasks for quicklook analysis: d2 displays two consecutive images, and just needs a 3-digit number as argument (eg d2 353 displays r935353[1] and r935354[1] to frames 1 and 2), and the famous seeing script that converts a FWHM of a star trail into seeing in arcsec.
Logsheet archive: All logsheets - also those of previous runs - were photographed and will soon appear on a website near you. Backup programme: After several nights of 2arcsec seeing, it would really be nice to have a useful backup programme. We thought one up, but no doubt there are others. Switching filters is now a 10-minute procedure and so not a real limitation. The coordinates of the local group dwarfs are in the back of one of the green PNS folders now, together with a record of what we looked at. In bad conditions this programme could be continued.

Run #11: 13-21 February 2007



Run #12: 13-21 October 2007

This was a technical visit by ND and Erik de Wit of ElectronX design (edewit.at.electronx.nl) in which Controller-II was connected to the PN.S (including the H-aC) to test the (C++) program, serial link, and the controller itself. In the months preceding Erik had revampt the hardware, including the front panel, backplane, and the like. There is a more complete summary coming HERE...
Run #13: December 2007

Despite the unlucky number, this was a good observing run for the PN.S and for the H-aC which obtained its first-light image. Here is Arianna's summary:

     NGC891north                   6.4
 
     NGC891south                   3.8
 
     NGC 2681                      3.1
 
     NGC 3115                      2.6
 
     NGC 936                       0.5
 
     NGC 3489                      0.7
 
 
  PS Total effective hours observed for each galaxy can be found in PNS
  webpage now updated (see below).
 
 
  -WEATHER, SEEING AND TIME LOST
 
     NIGHT 1
 
     Focus and readout noise suboptimal, entire night photometric,
  seeing from 1.1 arcsec to 1.3 arcsec. Missed 2 hours
  because autoguide was broken (duty engineer called).
  
     NIGHT 2
 
     Entire night photometric, seeing from 1.3 arcsec to 2 arsec
  (changeable because of wind), still some problem with the giude star.
 
     NIGHT 3
 
     Entire night photometric, very bad seeing, from 1.2 arcsec to 3
  arcsec, we were not really able to focus.
 
     NIGHT 4
  
     Entire night photometric, seeing from 1.1 arcsec to 2.4 arcsec.
  Understood how to use manually Halpha camera (Halpha broad band filter
  used).
 
     NIGHT 5
 
     Entire night photometric but very bad seeing from 2.2 arcsec to
  incredible values, from midnight to 3 a.m. we decided to close the
  dome because seeing was too bad for guiding, in the while we did dome
  flats.
  
     NIGHT 6
 
     Lots of clouds (ineable to guide for two hours) after midnight but
  good seeing from 1.1 arcsec to 1.4 arcsec.
 
 
  -Halpha CAMERA
 
  Halpha camera has been tested. The third day Edo understood how to use
  it manually: UNIBLITZ shutter (shutter of the Halpha camera) works in the inverse way.
  The fourth night we used broad band filter and the last two nights the narrow one. Taken very
  nice images with different exposure time in order to map all the
  galaxy shape from the center to the border.
 
  -PNS WEB PAGE
 
  The PNS log file in the web page has been update (now is possible to
  find all the effective hours observed for each galaxy in all the
  runs!! Thanks Edo!) and the new procedure used for the new Halpha
  camera has been written (thanks Lodo!).




Run #14: 31 March - 9 April 2008

please enter / check your attendence here
DATE: 31/3 1/4 2/4 3/4 4/4 5/4 6/4 7/4 8/4 9/4
NRN YES YES YES              
AC YES YES YES YES YES YES YES YES YES YES
LC         YES YES YES YES YES YES
FdL   YES YES YES YES YES YES YES    
notes:
  • we need to install new executable for laptop
  • LC arriving from Tenerife with Binter, the 4th morning. Contact (germ. mobile): +49 17666671507


1st night - March 31st

Observers: A. Cortesi, N.R.Napolitano (+R. Corradi as supporter)

AB0 filter on
Halpha not working, CCD had no signal.

NGC 3384 observed:  1.5 hrs; seeing ~2"5; effective observed time ~0.2hrs
NGC 4278 observed:  5.0 hrs; seeing ~ 3"0; effective observed time ~0.5 hrs

2nd night - April 1st

Observers: A. Cortesi, N.R.Napolitano, F. De Lorenzi

AB0 filter on
Halpha narror band filter ON
Right arm too high temperature: not working.
Observations taken only with Halpha Narrow and Left arm.
ToO observed (GRB)

NGC 3384 Left arm   observed 0.5 hrs ; seeing 1"3; effective ~0.3hrs
NGC 3384 HaN         observed 0.6 hrs ; seeing 1"3; effective ~0.4hrs

NGC 3115 Left arm   observed 0.5 hrs; seeing ~2"6; effective ~0.1 hrs
NGC 3115 HaN         observed 0.4 hrs; seeing ~2"6; effectve ~0.1 hrs

NGC 3489 Left arm   observed 1 hrs; seeing ~ 2"6; effective ~0.2 hrs
NGC 3498 HaN         observed 0.8 hrs; seeing 2"6 ; effective ~ 0.1 hrs

NGC 4278 Left arm   observed 2.0 hrs; seeing 2"3; effective 0.4 hrs
NGC 4278 HaN         observed 1.1 hrs; seeing 2"3; effective 0.2 hrs


3rd night - April 2nd
Observers: A. Cortesi, N.R. Napolitano, F. De Lorenzi

AB0 filter on
Halpha narror band filter ON
at 1:00 UT, closed for clouds

NGC 3384 L+R   observed 2.0 hrs; seeing ~1"5; effective ~0.9 hrs
NGC 3384 HaN observed 0.85 hrs; seeing ~1"5; effective ~0.4 hrs

NGC 4278 L+R     observed 1.5 hrs; seeing ~ 1"7; effective ~0.5 hrs
NGC 4278 HaN     observed 0.9 hrs; seeing 1"7; effective ~0.3 hrs


4th night - April 3rd

Observers: A. Cortesi, N.R. Napolitano, F. De Lorenzi

AB0 filter on
Halpha narror band filter ON
Observations ended at 3:50 UT for humdity

NGC 3384 L+R     observed 2.5 hrs; seeing~2"7; effective ~0.3 hrs:
NGC 3384 HaN     observed 1.5 hrs; seeing~2"7; effective ~0.2 hrs:
NGC 4278 L+R     observed 2.0 hrs; seeing~3"0; effective ~0.2 hrs:
NGC 4278 HaN     observed 2.0 hrs; seeing~3"0; effective ~0.2 hrs:



5th night - April 4th

Observers: A. Cortesi, L. Coccato, F. De Lorenzi

NGC 4278 L+R     observed ~2.5 hrs; seeing~1"2; effective ~1.6 hrs:
NGC 4278 HaB     observed ~2.1 hrs; seeing~1"2; effective ~1.4 hrs:
NGC 5866 L+R     observed ~0.75 hrs; seeing~1"2; effective ~0.3 hrs:
NGC 5866 HaB     observed ~0.75 hrs; seeing~1"2; effective ~0.3 hrs:


6th night - April 5th

Observers: L. Coccato, F. De Lorenzi

We close at ~22:40 (UT) because of the humidity.
NGC 3384 L+R     observed 1.5 hrs; seeing~2"7; effective ~0.2 hrs:
NGC 3384 HaB     observed 0.8 hrs; seeing~2"7; effective ~0.1 hrs:
NGC 3384 HaN     observed 0.6 hrs; seeing~3"0; effective ~0.1 hrs:
"

6th night - April 6 -- 8th

Observers: L. Coccato (6-8 april), F. De Lorenzi (6 april); A. Cortesi (6-8 april)

Not observed for high humidity
6th night - April 6 -- 8th

Observers: L. Coccato, A. Cortesi

NGC 3384 L+R     observed 0.5 hrs; seeing~3"0; effective ~0.06 hrs:
NGC 3384 HaN     observed 0.5 hrs; seeing~3"0; effective ~0.06 hrs:



last modified by L.Coccato apr 2008


Run #15: 27 - 30 Nov - 3 2008 (O8A/N10)

notes:
  • Using software 2.4 on this run.
  • Fresh gloves, and now dedicated box for A, B, AB, C
  • Observers LC ND AC SB (Samantha Rix supporting, replacing Romano)

Goal of run was bring several galaxies to completion

1st night - November 27
AB filter was in PN.S from previous run
PROBLEMS: the PNS had actually been mounted on Dec 26 but the Prime focus instrument still needed to be removed and/or balanced and so we had no profit from the early mount. On Dec 27 there was a minor road accident which delayed the crew. Andy Ridings was again struggling with serious earthing issues, and by 17:00 EEV12 was still noisy. Flats were missed entirely. We took turns to take dinner. At 20:00 we had still not started focussing the arms. Andy pushed on with the noise problem in a systematic way and we ended up with a very satisfactory L:4.1 R:3.3 H:3.9 with all 3 arms connected.
By this time NGC was lost for the night.
Next we discovered that no light was reaching the detectors. All connections seemed to be ok and shutter audibly moving. Turned out that focal station was wrong ("broken cassegrain") Telescope focus went well with the Focus routine.
  

results so far:


NGC 1161:

1.6 hours effective

1.5 hours in H-alpha narrow band

 NGC 2683: 1.1h effective  filter AB0

  NGC 2683: 1.1h effective  filter H-a narrow

  NGC 3384: 1.3 effective filter AB0

NGC 3384: 1.1 effective filter H-a narrow



2nd night - November 28 LC and ND went to focus instruments at 16:00 only to discover that the OII arms were in perfect focus with a flat pattern of spots of size 2.8 p/m 0.2.On the HaC (narrow) just as good, flat pattern at 3.0. On a suggestion of new support person Samantha Rix, SB has written a telescope focus script with same principle (best of a set of images) but much better in that the steps, including focus shift are automated by use of DAS scripts. NGC 584 targeted, so far 1"-1.2" seeing, no precipitation, and a beautiful sky. Seeing so good that we notice the difference when the telescope is refocussed from 98.0 to 98.05 NGC 584: 4h
RUN # 16 20-29 March 2009 Support astronomer: S. Rix
1st night, 20 Mar. Observers: Lodovico Coccato, Vincenzo Pota Telescope Operator: Norberto Gonzales
Filter: B0 - AB0 - HaN Note: On the PN.S. label it was written the filter AB0 was mounted. It was not true: filter B0 was in. Correction applied.
Bad seeing conditions, from a good 1.1 arcsec in the first 30 min, we increased 1.4...1.6...2.0...2.4 Dome closed at 0:30 for high humidity.
2nd night, 21 Mar. Observers: Lodovico Coccato, Vincenzo Pota Telescope Operator: Norberto Gonzales Filter: AB0 - HaN Bad seeing conditions. (minimum 1.3, maximum 2.4)
3rd night, 22 Mar. Observers: Lodovico Coccato, Vincenzo Pota Telescope Operator: Norberto Gonzales Filter: AB0 - HaN Good seeing conditions (0.9 - 1.2)
4rd night, 23 Mar. Observers: Lodovico Coccato, Nicola Napolitano, Vincenzo Pota Telescope Operator: Norberto Gonzales Filter: AB0 - HaB Bad seegin conditions (1.7 -- 3.8) Dome closed for humidity at 3:20 a.m. Night entirely lost. Good seeing conditions (0.9 - 1.2)
5th night, 24 Mar. Observers: Lodovico Coccato, Nicola Napolitano, Vincenzo Pota Telescope Operator: Antonio Garcia Dome closed for humidity
6th night, 25 Mar. Observers: Lodovico Coccato, Arianna Cortesi, Nicola Napolitano, Vincenzo Pota Telescope Operator: Antonio Garcia Dome closed Spanish science minister visiting the WHT.
7th night, 26 Mar. Observers: Arianna Cortesi, Nicola Napolitano, Vincenzo Pota Telescope Operator: Antonio Garcia Dome closed: wet, thunderstorms!
8th night, 27 Mar. Observers: Arianna Cortesi, Nicola Napolitano, Vincenzo Pota, Koen Kuijken Telescope Operator: Antonio Garcia Dome closed: wet, windy, freezing.
9th night, 28 Mar. Observers: Arianna Cortesi, Vincenzo Pota, Koen Kuijken Telescope Operator: Neil O'Mahony Clouds lifted in the afternoon, and we could observe. Seeing varied from 2 to 1.2 arcsec. Accumulated data on NGC 4278 and 4697. About 1 hour before twilight humidity rose again and we had to close up for half an hour. After the humidity recovered we observed half an hour on 5866 and a flux standard.
10th night, 29th March Observers: Arianna, Vincenzo, Koen Telescope operator: Neil Clear skies, seeing gradually improved from 2 to 1.2 arcsec before worsening at end of night. We lost the PNSL arm in the middle of the night - apparently an LED in the CCD dewar was on, saturating images in that arm (and affected the PNSR arm through internal reflections). Multiple attempts to switch controllers failed, so in the end we continued observing with H-alpha plus PNSR only. Accumulated data on NGC3115 (3 arms), 4278 (no L arm) and 5866 (ditto). END OF RUN

GENERAL COMMENTS (in progress...) The 3 CCDs readout noise are much more stable now. We experienced dificulties in focussing the AB0 filter. Rarm on the bottom right part of the image higher values of imexam-a-gaussian (~3.4). In general, focus on Rarm is worst than the Left. Situation have been (partially) solved the 4th night, after many and many tries. Nevertheless, the FWHM measured on the star trails is better in the Rarm than the Larm. On the 3rd night we measures seeing 0.9-1.0 on the Rarm and 1.0 - 1.1 on the Left. The first two nights we notices similar differences. We made an estimates of overheads based on observations of the 3rd night. We moved to the galaxy at 10:00p.m and observed it untill 6:00a.m. (8hrs) The total exposure time is 6.3 hrs (12 x 30min + 20min) Therefore the time lost (arcs, pointing, reading the CCD, focus of the telescope) is about 1.7hrs (21%). During this run we did some dome flats. Procedure on the PN.S webpage has been updated (the lamps are controlled form the control room now). Difference with the original procedure for dome flats is that now the shutter is used to control the exposure time, not the lamps themselves. So any shutter-related nonuniformity will be present in the dome flats. We took some test exposures of different duration to test for this effect. Tests with dome flats at different position angles proved that the pupil is not uniformly illuminated: a 180 degree rotation can lead to almost 50% changes in the ratio of the fluxes that go into each arm. Using all three lamps instead of the recommended 1 largely fixes this issue.
EXPOSURE TIMES (effective hours)
TARGET L/R HaN HaB NGC 2974 1.3 1.2 0.0 NGC 3489 3.0 2.2 0.5 NGC 4472N 4.4 4.4 NGC 4472S 4.0 3.6 0.3 NGC 5866 0.9 1.0 NGC 3115 0.6 0.4 NGC 4697 1.1 0.5 NGC 4278 1.6 3.0 (for NGC 4278 0.9h only on the Left arm)
Standard stars: BD+33 2642 SP0642+021 HZ44 SP1312+363


RUN # 17 11-17 Decmeber 2009 Support astronomer: S. Rix (?)
1st night, 11 Dec. Observers: Arianna Cortesi, Steven Bamford Telescope Operator: ?
Filter: AB0 -B0 - HaN Note: Changed filter from AB0 (that was in) to B0. Correction applied as in the webpage, the focus is good without any change.
Cloudy, rainy. We never opened, we went back at 3:30.
2nd night, 12 Dec. Observers: Arianna Cortesi, Steven Bamford Telescope Operator: ? Filter: B0 - HaN We never opened.
3rd night, 13 Dec. Observers: Arianna Cortesi, Steven Bamford Telescope Operator: ? Filter: B0 - HaN Humidity 100%, we never opened.
4rd night, 14 Dec. Observers: Arianna Cortesi, Steven Bamford Telescope Operator: ? Filter: B0 - HaN Dome closed for humidity 100%, living in a cloud. Trouble with PNSL , left arc images have degraded quality including ghosting aroud central slit. Suspected condansation somewhere. Discovered air flow across cryostat windows not working - pushing tubes into source box (with glass indicators ) restored flow.
5th night, 14 Dec. Observers: Arianna Cortesi, Steven Bamford Telescope Operator: ? Dome closed for humidity. We opened and closed twice without getting any good data.
6th night, 15 Dec. Observers: Arianna Cortesi, Steven Bamford Telescope Operator: ? Dome closed High wind!!

GENERAL COMMENTS (in progress...) 66 potential hours 3 minutes eff times observe!!!

RUN # 18 10-19 march 2010 Support astronomer: S. Rix
Observers: A. Cortesi, L. Coccato, N. Napolitano (15-18 March)

Summary of observed targets:
EFFECTIVE EXP. TIMES  (Texp/seeing2)

M101
   R band      :   0.4 hrs

NGC 2403
   L/R  AB6   :  5.4  hrs
   Halpha Narrow: 4.2  hrs
   R band:  2.3  hrs


NGC 2681
   L/R AB0 :    0.5 hrs
   Rband    :  0.5 hrs


NGC 2786
   L/R AB6    0.2 hrs
   L/R AB0    1.6 hrs
   Ha Narrow  1.2 hrs
   R band     0.5 hrs


NGC 3115
   L/R AB0     1.0 hrs
   Ha Narrow:  1.0 hrs


NGC 3377
   R band     0.08 hrs

NGC 3489
   L/R AB0        0.9 hrs
   L/R AB6        0.6
   Halpha Narrow  1.5 hrs
  NGC 4472 East
   L/R AB0     4.3 hrs
   Ha Narrow   2.9 hrs
   R band      1.3 hrs

NGC 4472 West    total L/R:      4.8 hrs
   Halpha Narrow:  2.7 hrs
   R band:         1.6 hrs

NGC 4472 North (extra field)
   total L/R:       4.2 hrs
   Halpha Narrow:   4.2 hrs

NGC 4472 South (extra field)
   total L/R:      4.9 hrs
   Halpha Narrow:  3.3 hrs
   R band     :    1.5 hrs

NGC 5866
    L/R AB0:        5.1 hrs
    Halpha Narrow:  3.6 hrs
    R band :        1.5 hrs 


Details for every night:

******************
 PNS RUN #18
 10-19 March 2010
*******************

Observers: L. Coccato, A. Cortesi, N. Napolitano


General remarks
----------------

New ccd quality control software, by S. Rix.

We have now a script that monitors the measures the out noise on
overscan regions on the images taken during the night.  It is very
useful to monitor the variation of R.O.N. during the night.
To activate it:
   source aliases
   ccdqcmon start  (stop to terminate it)


The ING staff was able to lower down the RON in all the detectors (it
ranges form 2.9 to 3.3 roughly).


Nights report:
**** Night 1: 10 March 2010           ****
**** Observers: L.Coccato; A.Cortesi  ****
B
ad seeing at the beginning of the night (~2)
Clear sky (till 00:50).
Left/Right filter change took ~1 hour (AB6--> AB0).
Dome closed at 00:50 for high humidity.

Targets observed:


Galaxy      Set Up    Texp  seeing   Teff(Texp/seeing2)
                             arcsec  
========================================================
NGC 2403    AB6 L/R   1800   >2       <450
           HaN       1820   >2       <455

NGC 2768    AB6 L/R   3600   ~2       ~900
           HaN       3200   ~2       ~800

NGC 4472S2  AB0 L/R   1800   ~2       ~450
           HaN       1800   ~2       ~450

NGC 3115    AB0  L/R  ~900   ~2       ~225
           R-band    ~800   ~2       ~200




**** Night 2: 11 March 2010           ****
**** Observers: L.Coccato; A.Cortesi  ****

Clear sky, with some veils showing up and going away.
Bad seeing at the beginning of the night, but it improved.

Technical problem with the telescope focus at the beginning of the
night. Solved.

Total effective (only second night):
   NGC 2403       ~1903 L/R;   1903 HaN   (~0.5 hrs)
   NGC 2768        ~918 L/R;    918 HaN   (~0.3 hrs)
   NGC 4472 S2   ~11067 L/R;  11000 HaN   (~3.1 hrs)
   NGC 5866        2885 L/R    2885 HaN   (~0.8 hrs)


**** Night 3: 12 March 2010           ****
**** Observers: L.Coccato; A.Cortesi  ****


Very good seeing at the beginning of the night (<1)
Some veils passing by(all the night long).
Left/Right filter change took ~11 minutes (AB6--> AB0)  :) .


Targets observed:


Galaxy      Set Up    Texp  seeing   Teff(Texp/seeing2)
========================================================
NGC 2403    AB6 L/R   3*1800   1.2       =5953.1
HaN       3*1800   1.2       =5953.1

NGC 2768    AB0 L/R   2*1800   ~1.1       =2737.6
HaN       2*1800   ~1.1       =2737.6

NGC 4472W  AB0 L/R    10*1800   ~1.4       ~9074 (= 2 h 30 min)
HaN       10*1800   ~1.4       ~9074

NGC 5866    AB0  L/R  ~2*1800   ~1.0       ~3600
R-band    ~2*1800   ~1.0       ~3600



*** Night 4: 13 March 2010           ****
**** Observers: L.Coccato; A.Cortesi  ****


Clear sky, seeing ranging from 0.9 to 1.3.
Our average seeing was around 1.1, while the Robo-dimm
was measuring 0.5-0.6.


Targets observed:


Galaxy      Set Up    Texp  seeing   Teff(Texp/seeing2)
                     (sec)           (sec)
========================================================
NGC 2403    AB6 L/R   1800x2  0.9    ~2222x2   (1.2 hrs)
           R-band    3060    0.9     3778

NGC 2768    AB0 L/R   1800    0.9    ~2222     (0.6 hrs)
           R-band    1480    0.9    ~1640

NGC 3489    AB0 L/R   1800    1.2    ~1250     (0.3 hrs)
           R-band    1560    1.2    ~1083

NGC 4472 S2 AB0 L/R   1800x5 ~1.2    ~6355     (1.7 hrs)            
         R-band    7400   ~1.2    ~5100

NGC 4472 W  AB0 L/R   1800x5 ~1.1    ~1488x5   (2.1 hrs)            
         R-band    7400   ~1.1     6115

NGC 5866    AB0 L/R   1800x2 ~1.0    1800x2    (1hrs)
           R-band



(Plus 3/4 spectrophotometric standard stars )

PS:

NGC 2768    Completed
NGC 4472 S2 Completed  (total 4.8 hrs)
NGC 4472 W  Completed  (total 4.6 hrs)



**** Night 5: 14 March 2010                          ****
**** Observers: L.Coccato; A.Cortesi, N. Napolitano  ****


Few clouds, seeing ranging from 1.0 to 1.6.
Our average seeing was around 1.4, while the Robo-dimm
was measuring 0.8.

On average, the night was not as good as the previous one.  Still not
clear to me the source of the difference between the seeing we
measure, and the seeing measured by the Robo-dimm.



Targets observed:


Galaxy      Set Up    Texp  seeing   Teff(Texp/seeing2)
                     (sec)           (sec)
========================================================
NGC 2403    AB6 L/R   1800x2  1.1    ~1488x2   (0.8 hrs)
           HaN       1800x2  1.1     1488x2

NGC 3489    AB0 L/R   1800x2  1.35    ~988x2   (0.54 hrs)
           HaN       1800x2  1.35    ~988x2

NGC 4472 N2 AB0 L/R   1800x10 1.4     ~918x10  (2.55 hrs) 
            HaN       1800x10 1.4     ~918x10  (2.55 hrs)

NGC 5866    AB0 L/R   2700    1.3      1600    (0.44hrs)
            HaN       2700    1.3      1600

(Plus 3/4 spectrophotometric standard stars )



**** Night 6th: 15 March 2010 6th night
**** Observers:  A.Cortesi, N. Napolitano


Weather conditions: Clouds, very high humidity at the beginning of
night, that overcome the 90% limit at 22:00, therefore we had to
close.  We opened again at 2:00, the humidity from then on dropped
off.  The seeing was from 1.4 to 2.0.

Seeing: The Robo-dimm was always not working apart at the end of the
night when our seeing was 1.4 and it was showing 0.8.

On average, the night was not as good as the previous one. Because of
the poor seeing, we switched to the backup program and we put in the R
filter in during the time the dome was closed. When we were on sky
again, since the seeing was below 1.8, we then kept observing the
galaxies listed in the main program.

Targets observed: 

Galaxy Set Up Texp seeing Teff(Texp/seeing2) (sec) (sec) 

======================================================== 

NGC 2403   AB6 L/R 1800x2 1/2 2 ~1200 (0.3 hrs) 
           HaN 1800x2 1/2 2 1200 

NGC 4472 E AB0 L/R 1800x7 1.4 ~918x10 (1.8 hrs) 
          HaB various 1.4 ~918x10     (<1.8 hrs) 

NGC 5866   AB0 L/R 1800 1.4 918 (0.44hrs) 
           HaB various 1.4 918 

(Plus 1 standard star ) 

 

**** Night 7: 17 March 2010                          ****
**** Observers: L.Coccato; A.Cortesi, N. Napolitano  ****


Cloudy, especially at the end of the night there was really poor
transparency, seeing ranging from 1.3 to 1.9. At the beginning of the
night, we performed a short exposure in Halpha in order to be compared
its image quality with with the robodim measurements. We measured a
seeing of 1.2 on an Halpha exposure of 60 sec, while the robodim was
showing 1.1. The discrepancy is not so large.

We also noticed that the L arm, that has the better focus in the
spectrograph, has a worse seeing measured on sky than the R arm and
the Halpha.

Targets observed:


 Galaxy      Set Up    Texp  seeing   Teff(Texp/seeing2)
                       (sec) 
========================================================
 NGC 2403    AB6 L/R   1800x2  1.4     (0.6 hrs)
             HaN       1800x2  1.4


 NGC 4472 E AB0 L/R   1800x10  1.4     (2.4 hrs)
             HaN      1800x10  1.6     (2.4 hrs)

 NGC 5866    AB0 L/R   3000    1.8      (0.3 hrs)
             HaN       3000    1.8      (0.3 hrs)

 (Plus 1 spectrophotometric standard stars at the beginning of the night) 



 ***** Report 8th night
 ***** Observers: L.Coccato, A.Cortesi, N. Napolitano

      
Weather conditions: all the night was characterized by the passage
of clouds that were either hiding the object or the guiding star, so we
changed object frequently trying to find gaps in the sky.  The seeing was
from 1.1 to 1.7. The Robo-dimm was always not working.

We extracted the a preliminary catalogue for NGC3115 and estimated a
maximum number of PNe of 130, therefore we kept observing NGC3115 at
the nominal position as there is no additional time to complete the
two external pointings.

     Targets observed:

Galaxy    Set Up    Texp   seeing   Teff(Texp/seeing2)

========================================================

NGC 2403   AB6 L/R  1800x1    1.4   0.26 hrs
           HaN      1800      1.4   0.26 hrs

NGC 3115   AB0 L/R  4600      1.2   0.9  hrs
           HaN      4600      1.2   0.9 hrs

NGC 4472N2  AB0 L/R  1800x4    1.5    0.9 hrs
            HaN      1800x4    1.5    0.9 hrs

 NGC 5866   AB0 L/R  1800*3    1.4    0.9 hrs
            HaN      1800*3    1.4    0.9 hrs




**** Night 9: 18 March 2010                          ****
**** Observers: L.Coccato; A.Cortesi, N. Napolitano  ****


Not a good night from the observing conditions: a strong wind made
very it difficult to track on the targets in the wind direction, and
we were forced to point the telescope in a different part of the sky
. We also had clouds and bad seeing.  The dome was closed for ~3 hours
because the wind speed exceeded the limits.

The seeing was from 3.3 (beginning of the night) down to 1.4 (end of
night). The average seeing was ~1.8. Robodimm was not operative
(wind). At the INT, observers measured bad seeing too.

Targets observed:


Galaxy      Set Up   Effective (exptime/seeing2)

========================================================
NGC 2681    AB0 L/R   0.5 hrs (the mean seeing was >2)
            Rband     0.5 hrs


NGC 4472 N2 AB0 L/R   0.7 hrs (the mean seeing was ~1.8)
            HaN       0.7 hrs


NGC 5866    AB0 L/R   0.2 hrs (the mean seeing was ~1.5)
            Rband     0.2 hrs


**** PNS Run 18; P10; N3                             ****
**** Night 10: 14 March 2010                          ****
**** Observers: L.Coccato; A.Cortesi, N. Napolitano  ****


Not a good night for the weather conditions: it was very cloudy, windy
and the dome closed for 5 hours. It was rather difficult to find a
guide star.

Average seeing was 2.2 arcsec, and it became even worst at the end of
the night.

We decided therefore to swap tp the backup program and do broad band
photometry. The targets are listed below:

NGC 3377 (It is an S0, and the catalog is already out. It can be included
          in Arianna's thesis).

M101     (the only bright target visible trough clouds).

NGC 5866 (observed at the end of night, when the sky cleared
         up a bit and it was possible to find a guide star).

Targets observed:


Galaxy      Set Up   Texp       seeing   Teff(Texp/seeing2)

============================================================
NGC 4472 E   AB6      1.4  hrs    2.2      0.3 hrs
             HaN      1.4  hrs    2.2      0.3 hrs

NGC 3377     R band   0.45 hrs    2.2

M 101        Rband    0.40 hrs    2.2

NGC 5866     AB0 L/R   1.5 hrs    2.7      0.20 hrs
             R band    1.3 hrs    2.7      0.18 hrs 



******************************************
 PNS RUN #19
 30 Dec 2010 6 Jan 2011
******************************************

Observers: 
M. Arnaboldi (nights: 4-6); L. Coccato (nights: 1-8), A. Cortesi(nights: 1-4)

** Night 1
100% time lost due to technical reasons and clouds.

Major failure on PNS left and righ controller, probably due to the
hardware migration to linux machines, which does not support our SUN
scripts at 100% (integration never stops, manual finish commands do not work).


We proceeded with back up program: Halpha imaging of sample galaxies
(+ one ToO: ACAM sloanR images of the asteroid Scheila)



** Nights 2 - 6
Problem of previous nigth solved by the ING staff.
During these nights we experimented few controller failures, which 
were solved by DASRESET. Only few exposures lost in total.
Nights 2-3 the Right arm RON was ~6 ADUs.

We got various seeing values, ranging from 1.0 to 2 arcsec, and the sky 
was clear (almost) all the time.
3rd and 4th nights only were cloudy.

** Night 7 the seeing was very good, ranging from 0.9 to 1.4, sky clear.

** Night 8 very good seeing (~1.0 - 1.1). 2 hours lost because of
2 override observing requests (GRB 101225 and GRB 110106B)


--------------
Run summary:
--------------


B-filter targets:

Targets observed       Total exposure time      Exposure time scaled to 1 arcsec
                           hrs                         hrs
=============================================================================
NGC 584  
          L/R B0       9.5(+1.5 only Red arm)    6.3(+1.3 only Red arm)         
          Ha Narrow       8.0                       5.4
	    Ha Broad        2.6                       1.9
-----------------------------------------------------------------------------
NGC 2549               
          L/R B6         17.4                       8.9 
          Ha Narrow      13.45                      8.2
          Ha Broad        6.3                       3.3           
-----------------------------------------------------------------------------
NGC 2974
          L/R B0       4.25(+2.0 only left arm)   2.1 (+0.9 only Left arm)
          Ha Narrow       6.25                      3.0
          Ha Broad        0.0                       0.0          
------------------------------------------------------------------------------


AB-filter targets:

Targets observed       Total exposure time      Exposure time scaled to 1 arcsec
                           hrs                         hrs
=============================================================================
NGC  936
          L/R AB0        10.0                          7.8
          Ha Narrow       7.5                          5.7
          Ha Broad        2.2                          1.8
------------------------------------------------------------------------------
NGC 2403  
          L/R AB6         5.0                          3.3
          Ha Narrow       5.0                          3.3
          Ha Broad        0.0                          0.0 
------------------------------------------------------------------------------
NGC 2681
          L/R AB0         7.5                          6.0
          Ha Narrow       4.5                          4.0
          Ha Broad        2.7                          1.9
------------------------------------------------------------------------------
NGC 3115
          L/R AB0         6.75                         5.6
	   Ha Narrow       3.25                         2.8
	   Ha Broad        3.2                          2.5
------------------------------------------------------------------------------

Considering the previous runs, we can consider the observations of 
NGC 584, NGC 2549, NGC 2681, NGC 3311 completed (i.e. at least 10 equivalent hours
at 1 arcsec seeing).



******************************************
 PNS RUN #30
 01-10 March 2011; Observers A. Cortesi, L. Coccato
******************************************

### Night 1 
01 March 2011

Photometric night, seeing ranging from 1.1 to 1.4, with some peaks at 1.7 1.8.

Targets observed:
Targets observed       Total exposure time      Exposure time scaled to 1 arcsec
                           hrs                         hrs
=============================================================================
NGC 2683  
          L/R AB0         1.5                       1.0
	  Ha Broad        1.2                       0.9
-----------------------------------------------------------------------------
NGC 3489  
          L/R AB0         6.0                       3.0
	  Ha Broad        5.3                       3.1
-----------------------------------------------------------------------------
NGC 4278  
          L/R AB0         1.5                       0.7
	  Ha Broad        1.3                       0.7
-----------------------------------------------------------------------------

### Night 2 
02 March 2011

Veils, seeing ranging from 1.7 to 1.2

Targets observed:
Targets observed       Total exposure time      Exposure time scaled to 1 arcsec
                           hrs                         hrs
=============================================================================
NGC 2683  
          L/R AB0         1.5                       0.6
	  Ha Broad        1.35                      0.5
-----------------------------------------------------------------------------
NGC 3384  
          L/R AB0         5.5                       2.2
	  Ha Broad        5.1                       2.0
-----------------------------------------------------------------------------
NGC 4278  
          L/R AB0         2.0                       1.2
	  Ha Broad        1.8                       1.1
-----------------------------------------------------------------------------

### Nights 3-8
03-08 March 2011

Dome close: 100% Humidity 

#### Night 9
4 hrs lost for 100% humidity. Rest of the night thick clouds and bad seeing (1.7 - 2.5)
I would not use these images, but I put them in the list for the records.

Targets observed       Total exposure time      Exposure time scaled to 1 arcsec
                           hrs                         hrs
=============================================================================
NGC 2974
   L/R  B0                 1.5                         0.3
   Ha Narrow               1.5                         0.3
-----------------------------------------------------------------------------
NGC 4473
   L/R  B0                 3.0                         1.0
   Ha Narrow               3.0                         1.0
==============================================================================

### Night 10 - END OF RUN

100% Humidity.


******************************************
 PNS RUN #31
 14-23 March 2012; 
 Observers: A. Chies-Santos, L. Coccato, N. Napolitano
******************************************

### Night 1
14 April 2012
Observers: L. Coccato, N. Napolitano
Telescope Operator: Fiona Riddick
Support Astronomer: Marie Hrudkova

Sky conditions: clear, photometric
Seeing: variable (1.2 - 1.8)

Targets observed      Total exposure time      Exposure time scaled to 1''
                           hrs                         hrs
=============================================================================
NGC 4339
   L/R  B6                 5.0                         2.6
   Ha Narrow               5.0                         2.6
-----------------------------------------------------------------------------
NGC 5866
   L/R  AB0                2.0                         1.0
   Ha Narrow               2.0                         1.0
==============================================================================


Standard stars observed: 
    Hilt 600 (Mv ~11mag); Exposure: 30s (L/R-B6); 10s (HaN) 
    SP1550+330 (Mv ~11mag); Exposure: 30s (L/R-AB0); 5s (HaN) 


Technical details:
Offset to switch filter B6 --> AB0 (+ is clock-wise)
  L1  +0.55
  L2  +0.50
  L3  +0.65

  R1  +0.45
  R2  +0.45
  R3  +0.60


### Night 2
15 April 2012
Observers: L. Coccato, N. Napolitano
Telescope Operator: Fiona Riddick
Support Astronomer: Marie Hrudkova

Sky conditions: clear, photometric
Seeing: 
    beginning of the night: 1.2 -- 1.3
    middle of the night:    2.0 -- 1.7 
    end of night:           1.5 -- 1.3

Targets observed      Total exposure time      Exposure time scaled to 1''
                           hrs                         hrs
=============================================================================
NGC 4339             
   L/R  B6                1.5                           0.9                 
   Ha Narrow              1.5                           0.9              
-----------------------------------------------------------------------------
NGC 4742                  
   L/R  B0                2.5                           0.7
   Ha Narrow              2.5                           0.7
-----------------------------------------------------------------------------
NGC 5866
   L/R  AB0               2.5                           1.2
   Ha Narrow              2.5                           1.2
-----------------------------------------------------------------------------



Standard stars observed: 
    Hilt 600 (Mv ~11mag); Exposure: 30s (L/R-B6); 10s (HaN) 

    SP1550+330 (Mv ~11mag); Exposure: 30s (L/R-AB0); 5s (HaN) 



### Night 3
16 April 2012
Observers: L. Coccato, N. Napolitano
Telescope Operator: Fiona Riddick
Support Astronomer: Marie Hrudkova

Sky conditions: veils,  strong wind
Bad seeing 1.5 -- 2.2 


Targets observed      Total exposure time      Exposure time scaled to 1''
                           hrs                         hrs
=============================================================================
NGC 2974         
   L/R  B0                1.5                          0.57
   Ha Narrow              1.5                          0.57
-----------------------------------------------------------------------------
NGC 4742        
   L/R  B0                2.5                          0.63
   Ha Narrow              2.5                          0.63
-----------------------------------------------------------------------------
NGC 5866                  
   L/R                    2.0                          0.62
   Ha Narrow              2.0                          0.62
-----------------------------------------------------------------------------




Standard stars observed: 
    Hilt 600 (Mv ~11mag); Exposure: 30s (L/R-B6); 5s (HaN) 



### Night 4
17 April 2012
Observers: L. Coccato, N. Napolitano
Telescope Operator: Fiona Riddick
Support Astronomer: Marie Hrudkova

Sky conditions: veils, wind
  Seeing: 
    beginning of the night: ~2.0 than it got ~4.0
 
Time lost because of seeing 6hrs

Targets observed      Total exposure time      Exposure time scaled to 1''
                           hrs                         hrs
=============================================================================
NGC 2974         
   L/R  B0                1.5                          0.52
   Ha Narrow              1.0                          0.40
-----------------------------------------------------------------------------
NGC 4742        
   L/R  B0                1.5                          0.30
   Ha Narrow              1.5                          0.30
-----------------------------------------------------------------------------



Deep R-band imaging with Halpha Broad (seeing 2--4 arcsec):

  NGC 3608/3607   10x5 min 
  NGC 4339         7x5 min + 3x1 min
  NGC 5866        10x5 min 





### Night 5
18 April 2012
Observers: L. Coccato, N. Napolitano
Telescope Operator: Fiona Riddick
Support Astronomer: Marie Hrudkova

Sky conditions: calima, veils
  Seeing: 
    till midnight variable between  ~1.3 and 1.5
    from midnight ~1.1

Targets observed      Total exposure time      Exposure time scaled to 1''
                           hrs                         hrs
=============================================================================
NGC 2974         
   L/R  B0                0.5                          0.3
   Ha Broad               0.4                          0.3
-----------------------------------------------------------------------------
NGC 4365
   L/R  B6                5.5                          3.3
   Ha Narrow              5.5                          3.3
-----------------------------------------------------------------------------
NGC 5866                  
   L/R  AB0               1.5                          1.3
   Ha Narrow              1.5                          1.3
-----------------------------------------------------------------------------


### Night 6
19 April 2012
Observers: L. Coccato, A. Chies-Santos
Telescope Operator: Fiona Riddick
Support Astronomer: Raine Karjalainen

Sky conditions: few veils
  Seeing: ~1.0


Targets observed      Total exposure time     Exposure time scaled to 1''
                           hrs                        hrs
=============================================================================
NGC 4365 
   L/R  B6                 6.5                        6.4
   Ha Narrow               6.0                        6.0
-----------------------------------------------------------------------------
NGC 5866                  
   L/R  AB0                1.5                        1.2
   Ha Narrow               1.5                        1.2
-----------------------------------------------------------------------------



In two images obtained with the Ha camera of the PN.S, the stars are
elongated, as if the guide was lost or the telescope "jumped" on two
positions (yes Nicola, the problem occurred again tonight). This
feature is not observed in the Left and Right spectra, which are
simultaneous to the Halpha exposure. We therefore exclude a problem
related to the guiding or the telescope tracking.

The problem occurred for the same galaxy, at about the same time of the
night, so we think it is related to the position the telescope is
pointing.

The images are: r1782011.fit (20120418) and r1782195.fit (20120419)

We have two hypothesis:

1) The Halpha arm of the PNS is subject to flexures, which happen at a
   given coordinates. Maybe a loose screw.

2) The dichroic or the Halpha filters have some flexures, which happen
   at given coordinates on the sky. A flexure in the dichroic in fact
   will cause a change in the optical path of the Halpha camera, not
   of the OIII cameras (the dichroic is transparent to the OIII
   wavelengths).


Today I'll send a fault report, but I would prefer to have this problem
investigated **after** the end of the run. In fact, a part for this
inconvenient that happened only twice, the instrumentation is behaving
very well and we prefer not to touch it within the observing run.




### Night 7
20 April 2012
Observers: L. Coccato, A. Chies-Santos
Telescope Operator: Fiona Riddick
Support Astronomer: Raine Karjalainen

Sky conditions: clear
  Seeing:  1.0 -- 1.2, occasional peaks to 1.3-1.4 middle of the night.
 
  Seeing in Halpha 0.9 -- 1.0



=============================================================================
Targets observed      Total exposure time     Exposure time scaled to 1''
                           hrs                        hrs
=============================================================================
NGC 4365 
   L/R  B6                 1.0                        0.9
   Ha Broad                1.0                        0.9
-----------------------------------------------------------------------------
NGC 4742 
   L/R  B6                 4.0                        3.1
   Ha Narrow               1.5                        1.1
   Ha Broad                2.5                        2.0
-----------------------------------------------------------------------------
NGC 5363 
   L/R  AB0                1.5                        1.1
   Ha Narrow               1.5                        1.1          
-----------------------------------------------------------------------------
NGC 5866 
   L/R  AB0                1.0                        1.0
   Ha Narrow               1.0                        1.0
-----------------------------------------------------------------------------




### Night 8
21 April 2012
Observers: L. Coccato, A. Chies-Santos
Telescope Operator: Karl Kolle
Support Astronomer: Raine Karjalainen

Sky conditions: clear, strong wind, seeing ~ 1.8 - 2.2


=============================================================================
Targets observed      Total exposure time     Exposure time scaled to 1''
                           hrs                        hrs
=============================================================================
NGC 2974         
   L/R  B0                1.5                          0.4
   Ha Narrow              1.4                          0.4
-----------------------------------------------------------------------------
NGC 4742 
   L/R  B6                3.5                          1.0
   Ha Narrow              3.5                          1.0
-----------------------------------------------------------------------------
NGC 5363 
   L/R  AB0               1.5                          0.4
   Ha Narrow              1.5                          0.4
-----------------------------------------------------------------------------
NGC 5866 
   L/R  AB0               0.5                          0.1
   Ha Narrow              0.5                          0.1
-----------------------------------------------------------------------------





### Night 9
22 April 2012
Observers: L. Coccato, A. Chies-Santos
Telescope Operator: Karl Kolle
Support Astronomer: Raine Karjalainen

Sky conditions: clear
  Good seeing:
  beginning till 2:30:  0.8 -- 1.1 arcsec
  2:30 till 4:30:       1.2 -- 1.4 arcsec
  4:30 till end:        ~1.9  arcsec

Targets observed      Total exposure time      Exposure time scaled to 1''
                           hrs                         hrs
=============================================================================
NGC 4552            
   L/R  B6                  3.5                         3.6
   Ha Narrow                3.5                         3.6
---------------------------------------------------------------------------
NGC 4365 
   L/R  AB0                 1.5                         1.0
   Ha Narrow                1.5                         1.0
-----------------------------------------------------------------------------
NGC 5363  
   L/R  AB0                 1.5                         0.5
   Ha Narrow                1.5                         0.5
-----------------------------------------------------------------------------
NGC 5866 
   L/R  AB0                 0.25                        0.11
   Ha Narrow                0.25                        0.11
-----------------------------------------------------------------------------



### Night 10
23 April 2012
Observers: L. Coccato, A. Chies-Santos
Telescope Operator: Karl Kolle
Support Astronomer: Raine Karjalainen

Sky conditions: clear
  Good seeing:
  Beginning (till 21:30) ~2.0
  From 21:30 till 1:30    ~1.4
  From 01:00 till end: > 2.0

  Time Lost 3:30 hrs

Targets observed      Total exposure time      Exposure time scaled to 1''
                           hrs                         hrs
=============================================================================
NGC 4552            
   L/R  A0                  3.5                     1.4 
   Ha Broad                 3.0                     1.3
---------------------------------------------------------------------------
NGC 4365            
   L/R  AB0                 2.0                     0.5 
   Ha Broad                ~2.0                    ~0.5
---------------------------------------------------------------------------



Broad band imaging of NGC 5866  8x600 sec exposures at seeing ~2 or worst.
 



Tonight we mounted the A0 filter.
Offsets (- is counter-clockwise):

     AB0                       A0
   (absolute)     offset    (absolute)

L1  0.10          -0.10       0.00
L2  0.05          -0.10       0.95
L3  0.25          -0.10       0.15

R1  0.80          -0.10       0.70
R2  0.80          -0.10       0.70
R3  0.70          -0.10       0.60

PSF is about ~2.9 pixels (a-imexam, moffat profile)



####### RUN SUMMARY: 
Targets observed           Exp @ 1''     TOTAL
                             hrs         hrs
===================================================
NGC 2974         
   L/R  B0                    1.8         7.7
   Ha Narrow                  1.3         7.3
   Ha Broad                   0.3         0.9
---------------------------------------------------
NGC 4339             
   L/R  B6                    3.5        10.0         
   Ha Narrow                  3.5         3.5     
   Ha Broad                   0.0         0.0
---------------------------------------------------
NGC 4365
   L/R  B6                   10.7        10.7
   L/R  BAB0                  1.5	  1.5
   Ha Narrow                  9.7	  9.7
   Ha Broad                   1.2	  1.2
---------------------------------------------------
NGC 4552            
   L/R  B6                    3.6         8.3
   L/R  A0                    1.4         4.7
   Ha Narrow                  3.6         3.6
   Ha Broad                   1.3         1.3
---------------------------------------------------
NGC 4742            
   L/R  B0                    5.9        10.0
   Ha Narrow                  2.8         2.8
   Ha Broad                   3.8         3.8
---------------------------------------------------
NGC 5363                
   L/R  B6                    2.1         2.1
   Ha Narrow                  2.1         2.1
   Ha Broad                   0.0         0.0
---------------------------------------------------
NGC 5866
   L/R  AB0                   5.1        15.4
   Ha Narrow                  3.6        10.6
   Ha Broad                   2.6         2.6
==================================================


Question on NGC 4365: is B6 the Correct filter? Would AB0 have been better?



----------------
Notes on focus.
----------------

Note for the B filter.  
In the L/R arms, the seeing seems to have a spatial dependence, larger
on one side of the chip, smaller on the other. The difference is about
0.2 arcsec. This does not depend on the spectrograph focus procedure,
which shows nearly homogeneous focus all over the field. It is a bit
difficult to detect tilting, because the scatter of of the arc lines
from one hole to the other can be of about 0.3-0.4 pixels. In any
case, the psf of the mask spots are always better than the seeing.

Does it depend on the different lightpath between galaxy light (form
the infinite) and arc-lamp light (from ~1 meter)? The following test
suggests no.

We compare the FWHM of tungsten spectra (mask on, internal lamps) with
the mask of a sky (mask on, sky flat), and they seem the same.




Last modified by LC April 2012